# references.bib

### v2

@@ 4 A Battery of Agents
We provide implementations of three main agent types for running computational experiments in the Animal-AI Environment: Random Action Agent, Heuristic Agent, and Learner. Each of these showcase the versatility of the environment for running experiments in multiple disciplines. Implementation of these agents is available in the supplemental material (see Section Open Access Statement).
@@
When performing testing on artificial and biological agents, it is often useful to define chance performance on tests that have been developed. However, experiments might be very complex, meaning that chance performance is not easily defined analytically. We provide customizable random action agents to simulate chance performance. The number of steps before picking a new action can be defined, or drawn from one of several distributions, and the agent can be biased to pick certain actions over others, or to correlate future actions with actions it has previously taken.
@@ 4.1.1 Foraging Experiment
We studied four types of agents on the foraging task in Figure 11. To serve as a baseline for chance performance on this test, we used a random action agent. This agent selects one of 9 actions with equal probability. The number of steps to take performing that action is a positive number drawn from a clipped normal distribution with mean of 5 and a standard deviation of 1, rounded to the nearest integer. We also designed a heuristic agent, to provide an example of near-optimal performance on this task. This agent navigates towards good goals using raycasts (45 rays over the frontal 60 degrees). When it gets stuck (i.e., its velocity is near 0 in all dimensions), it moves forwards and left or forwards and right with a probability of 0.5 until its velocity increases again. We then trained two Deep Reinforcement Learning agents, Proximal Policy Optimisation (PPO) and Dreamer-v3, on the task, with 64$\times$64 pixel input for one million steps each. We used the XL variant of Dreamer-v3, which has 200M learnable parameters. We then evaluated all four agents on 100 episodes of the Foraging task. The training curves for PPO and Dreamer-v3, as well as the distribution of results for all four agents, are presented in Figure 14. Note that all training curves presented in this paper are smoothed using 100-episode partial moving averages. Dreamer-v3, PPO, and the Heuristic Agent performed similarly and above the level of chance defined by the random action agent. A Kruskall-Wallis Rank Sum Test indicates that there was a significant difference between the agents ($\chi^{2}$(3) = 318.16, p < 0.00001). Post-hoc Dunn Test statistics ([FSADunnTest]) indicate that all agents are significantly different from each other, except for Dreamer-v3 and the Heuristic Agent (see Appendix B.1).
@@ 4.1.3 The Animal-AI Testbed
PPO and Dreamer-v3 were trained on a cumulative curriculum consisting of 11 stages. The first stage contained one variant from each of 30 tasks in Food Retrieval (L1), totalling 30 configurations which the agent looped over for two million steps. The second stage contained one variant from each of the 60 tasks in L1 and Preferences (L2), totalling 60 configurations. The pattern repeats until the tenth stage, containing one variant from each of the 300 tasks in L1-L10. The agents were trained for two million steps on the first ten stages. They were then trained for a further five million steps on the last stage again (300 configurations from tasks L1-10), to ensure adequate experience with the all the tasks in the Testbed. Agents were therefore trained for a total of 25 million steps.
@@ 5 Conclusion
Contributions & Acknowledgements
Project Conception: L. G. Cheke, J. Hernández-Orallo, M. Crosby, M. Halina, K. Voudouris
Unity Development: M. Crosby, J. Holmes, I. Alhas, B. Slater
Python Development: M. Crosby, I. Alhas, W. Schellaert, K. Voudouris, B. Slater, M. G. Mecattaf
Agent Design: M. Crosby, W. Schellaert, K. Voudouris, J. Burden, N. Chaubey, M. Patel, M. G. Mecattaf
Experiments: K. Voudouris, W. Schellaert, M. G. Mecattaf
Writing (Original Draft): K. Voudouris
Writing (Revisions): K. Voudouris, I. Alhas, J. Burden, J. Holmes, W. Schellaert, L. G. Cheke, J. Hernández-Orallo, M. Halina, N. Chaubey, N. Donnelly, B. Slater, M. G. Mecattaf
Senior Authorship: L. G. Cheke, J. Hernández-Orallo, M. Halina
The authors would also like to thank Benjamin Beyret, Murray Shanahan, and Danaja Rutar for their contribution to earlier versions of the Animal-AI Environment.
Open Access Statement
Further information, tutorials, scripts, and documentation for the project can be found here: https://github.com/Kinds-of-Intelligence-CFI/animal-ai. Scripts to replicate the experiments in this paper, along with all datasets and analysis scripts, can be found here: https://github.com/Kinds-of-Intelligence-CFI/aai3-paper-experiments. None of the experiments reported here were preregistered.
@@ Appendix A Sources of Maintenance Statuses of Other Environments
We defined maintenance status according to the last commit on the respective GitHub repository for each environment, as of 20th June 2024. Further information is presented in Table 3.
Table 3: The dates of the last commit in the GitHub history of each environment presented in Table 1. AAI3 is omitted as it is under constant active development across several repositories. These data are correct as of 2024-Sep-09.
@@
MuJoCo 2024-Sep-09 https://github.com/google-deepmind/mujoco/commits/main/
MineRL 2024-Aug-21 https://github.com/minerllabs/minerl/commits/dev/
@@
MiniWorld 2024-Sep-27 https://github.com/Farama-Foundation/Miniworld/commits/master/

### v3

@@ 4 A Battery of Agents
We provide implementations of three main agent types for running computational experiments in the Animal-AI Environment: Random Action Agent, Heuristic Agent, and Learner. Each of these showcase the versatility of the environment for running experiments in multiple disciplines. Implementation of these agents is available in the supplemental material (see Section Availability of Data and Materials).
@@
When performing testing on artificial and biological agents, it is often useful to define chance performance on tests that have been developed. However, experiments might be very complex, meaning that chance performance is not easily defined analytically. We provide customisable random action agents to simulate chance performance. The number of steps before picking a new action can be defined, or drawn from one of several distributions, and the agent can be biased to pick certain actions over others, or to correlate future actions with actions it has previously taken.
@@ 4.1.1 Foraging Experiment
We studied four types of agents on the foraging task in Figure 11. To serve as a baseline for chance performance on this test, we used a random action agent. This agent selects one of 9 actions with equal probability. The number of steps to take performing that action is a positive number drawn from a clipped normal distribution with mean of 5 and a standard deviation of 1, rounded to the nearest integer. We also designed a heuristic agent, to provide an example of near-optimal performance on this task. This agent navigates towards good goals using raycasts (45 rays over the frontal 60 degrees). When it gets stuck (i.e., its velocity is near 0 in all dimensions), it moves forwards and left or forwards and right with a probability of 0.5 until its velocity increases again. We then trained two deep reinforcement learning agents, Proximal Policy Optimisation (PPO) and Dreamer-v3, on the task, with 64$\times$64 pixel input for one million steps each. We used the XL variant of Dreamer-v3, which has 200M learnable parameters. We then evaluated all four agents on 100 episodes of the Foraging task. The training curves for PPO and Dreamer-v3, as well as the distribution of results for all four agents, are presented in Figure 14. Note that all training curves presented in this paper are smoothed using 100-episode partial moving averages. Dreamer-v3, PPO, and the Heuristic Agent performed similarly and above the level of chance defined by the random action agent, evidence of efficient foraging strategies that approximately maximise reward. A Kruskall-Wallis Rank Sum Test indicates that there was a significant difference between the agents ($\chi^{2}$(3) = 318.16, p < 0.00001). Post-hoc Dunn Test statistics ([FSADunnTest]) indicate that all agents are significantly different from each other, except for Dreamer-v3 and the Heuristic Agent (see Appendix B.1).
@@ 4.1.3 The Animal-AI Testbed
PPO and Dreamer-v3 were trained on a cumulative curriculum consisting of 11 stages. The first stage contained one variant from each of 30 tasks in Food Retrieval (L1), totalling 30 configurations which the agent looped over for two million steps. The second stage contained one variant from each of the 60 tasks in L1 and Preferences (L2), totalling 60 configurations. The pattern repeats until the tenth stage, containing one variant from each of the 300 tasks in L1-L10. The agents were trained for two million steps on the first ten stages. They were then trained for a further five million steps on the last stage again (300 configurations from tasks L1-10), to ensure adequate experience with all the tasks in the Testbed. Agents were therefore trained for a total of 25 million steps.
@@ 5 Conclusion
Ethics Approval
Ethical approval is not required because no experiments were performed with humans or animals. All data from human children were from an openly accessible repository (https://osf.io/g8u26/) associated with an independent study that received ethical approval.
Consent to Participate
No research was conducted with human subjects. All data from human children were from an openly accessible repository (https://osf.io/g8u26/) associated with an independent study that received ethical approval.
Consent for Publication
No research was conducted with human subjects. All data from human children were from an openly accessible repository (https://osf.io/g8u26/) associated with an independent study that received ethical approval.
Code Availability
Further information, tutorials, scripts, and documentation for the project can be found here: https://github.com/Kinds-of-Intelligence-CFI/animal-ai. Scripts to replicate the experiments and analyses in this paper can be found here: https://github.com/Kinds-of-Intelligence-CFI/aai3-paper-experiments.
Author Contributions
K. Voudouris: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Resources, Data Curation, Writing (Original Draft), Writing (Review & Editing), Visualization, Project Administration. I. Alhas: Software, Validation, Data Curation, Writing (Review & Editing). W. Schellaert: Software, Validation, Investigation, Data Curation, Writing (Review & Editing). M. G. Mecattaf: Formal Analysis, Software, Validation, Investigation, Data Curation, Visualization, Writing (Review & Editing). B. Slater: Software, Validation, Writing (Review & Editing). M. Crosby: Conceptualization, Formal Analysis, Software, Writing (Review & Editing). J. Holmes: Software, Writing (Review & Editing). J. Burden: Formal Analysis, Writing (Review & Editing). N. Chaubey: Formal Analysis, Writing (Review & Editing). N. Donnelly: Data Curation, Writing (Review & Editing). M. Patel: Software, Data Curation, Writing (Review & Editing). M. Halina: Conceptualization, Resources, Supervision, Funding Acquisition, Writing (Review & Editing). J. Hernández-Orallo: Conceptualization, Resources, Supervision, Funding Acquisition, Writing (Review & Editing). L. G. Cheke: Conceptualization, Methodology, Resources, Supervision, Funding Acquisition, Writing (Review & Editing).
Availability of Data and Materials
All datasets and materials for the experiments reported here can be found at https://github.com/Kinds-of-Intelligence-CFI/aai3-paper-experiments. None of the experiments reported here were preregistered.
@@ Appendix A Sources of Maintenance Statuses of Other Environments
We defined maintenance status according to the last commit on the respective GitHub repository for each environment, as of 17th January 2025. Further information is presented in Table 3.
Table 3: The dates of the last commit in the GitHub history of each environment presented in Table 1. AAI3 is omitted as it is under constant active development across several repositories. These data are correct as of 2025-Jan-17.
@@
MuJoCo 2025-Jan-16 https://github.com/google-deepmind/mujoco/commits/main/
MineRL 2024-Nov-26 https://github.com/minerllabs/minerl/commits/dev/
@@
MiniWorld 2025-Jan-12 https://github.com/Farama-Foundation/Miniworld/commits/master/